# There and Back Again: On the relation between Noise and Image Inversions in Diffusion Models

### v3

@@ 5 Conclusions
In this work, we analyze how errors introduced by the DDIM inversion algorithm influence the properties of latent image representations. Our findings reveal that the inverted latents exhibit systematic deviations from a Gaussian distribution, particularly in regions corresponding to smooth image surfaces. We trace this effect to the higher inversion error in these areas, which predominantly arises during the early noising steps. This discrepancy leads to practical limitations such as reduced diversity and lower text alignment to the conditioning prompts, or worse quality interpolation and editing in situations where latents are used as a starting point of the diffusion process. To mitigate this issue, we introduce a simple yet effective solution, where we decorrelate latents by replacing the initial steps of the inversion procedure with forward diffusion. Our approach significantly improves the quality of samples in the practical use cases of DDIM inversion.
@@
In the Appendix we first outline the limitations (A) of our experiments, discuss the broader impact (B) of the introduced method, and list the computational resources (C) used. We then describe, in detail, the DDIM approximation error (D), which effects we study in this work and the consequences it has across varying numbers of sampling steps (E).
Next, we describe our experimental procedures: measuring the angles of the noise–image–latent triangle (F), identifying plain-regions pixels (G) in the image, and computing the inversion error (H). In section I, we provide the full pseudocode for the proposed algorithm.
We demonstrate how the technique enhances image diversity and prompt alignment (J) and how it enables stochastic editing of real images (K). Finally, we compare Gaussian noise and DDIM latents in their mappings to images (L) and track how these relationships evolve during diffusion training (M). In section N, we include additional qualitative results for the experiments, especially when introduced method might fail, and how it performs in the task of image interpolation.
@@ Appendix A Limitations
In this work, we analyzed the relation between the random noise, image generations and their latent encodings obtained through DDIM Inversion. While our studies focused on DDIM approximation error from Song et al. [50], there exist other solvers and inversion methods, as described in section 2, bringing their own advantages and limitations. The error of DDIM inversion strongly depends on the number of steps with which it is performed. In particular, performing the process very granularly, e.g., using $T=1000$ steps, can result in strong suppression of the correlation. Nevertheless, the default number of steps we have chosen, i.e., 100, is, according to previous works [18, 12, 24], a practical choice as a good balance between the reconstruction error and the speed of the algorithm. In appendix E, we present that the latents exhibit correlations when using more sampling steps.
@@
The observations from our analytical experiments (correlations in table 2, angles in table 3, variance in table 4, interpolations in fig. 5) generalize well to all tested diffusion models, but are less evident in the LDM model trained on the CelebA-HQ images. We attribute this exemption to the specificity of the dataset on which the model was trained - photos with centered human faces, usually with uniform backgrounds. We believe that, unlike models trained on a larger number of concepts, the process of generating faces with uniform backgrounds is more stable and introduces little detail in subsequent steps, making the difference in approximation error for plain and non-plain areas less significant.
@@
As mentioned in section 2, the DDIM inversion error can be detrimental when using a small number of steps. Even though the solution proposed in this work (involving forward diffusion in first inversion steps) drastically removes correlations in latents and, thus, improves image interpolation and editing, it does not improve the numerical inversion error resulting from using small number of steps. In our experiments with 50 steps that are commonly used for edition, we show no significant drawbacks. However, in the extreme cases, using our method in even a single step, might result in the loss of information necessary for correct image reconstruction, hence it may be less preferred than standard DDIM inversion in some cases. In section N.1, we present failure cases for introduced solution.
Appendix B Broader impact
@@
As our work is mostly analytical, we do not provide new technologies that might have a significant societal impact. However, our solution for improving the accuracy of DDIM inversion has potential implications that extend beyond technical advancements in diffusion models. As our approach enables more precise image editing it could be misused to advance image manipulation techniques. The enhanced interpolation quality could make synthetic content more convincing and harder to detect.
Appendix C Compute resources
@@
Almost all of the experiments we perform are based on performing a sampling process using a diffusion model from noise, performing DDIM inversion and, possibly, image reconstruction from the latent, where each of these processes takes the same number of steps, hence the same number of GPU-hours on average. As our experiments differ in terms of number of sampling steps and images to generate, we provide average GPU time per one sampling step per batch (with $B$ denoting batch size) for each model as following: ADM-32 ($B=256$): 0.054s, ADM-64 ($B=128$): 0.093s, ADM-256 ($B=64$): 0.901s, LDM ($B=128$): 0.273s, DiT ($B=128$): 0.104s, IF ($B=64$): 0.609s. Note that some experiments, such as analyzing inversion approximation errors per step (fig. 4) or sampling from noise interpolations (fig. 5), involves performing the procedures several times.
Taking into consideration all the experiments described in the main text of this work, fully reproducing them takes roughly 95 GPU hours. However, considering the prototyping time, preliminary and failed experiments, as well as the fact that most of the experiments must be performed sequentially (e.g., inversion after image generation, image reconstruction after inversion), the overall execution time of the entire research project is multiple times longer.
Appendix D Approximation error in DDIM Inversion
Denoising Diffusion Probabilistic Models (DDPMs [16]) generate samples by reversing the forward diffusion process, modeled as a Markov Chain, where a clean image $x_{0}$ is progressively transformed to white Gaussian noise $x_{T}$ in $T$ diffusion steps. A partially noised image $x_{t}$, which serves as an intermediate object in this process, is expressed as $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$ where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, $\alpha_{t}=1-\beta_{t}$, $\epsilon_{t}\sim\mathcal{N}(0,\mathcal{I})$, and $\{\beta_{t}\}_{t=1}^{T}$ is a variance schedule, controlling how much of the noise is contained in the image at the specific step $t$.
@@ Appendix D Approximation error in DDIM Inversion
Song et al. [50] reformulate the diffusion process as a non-Markovian, which leads to a speed-up of the sampling process. While previously obtaining a less noisy image at $x_{t}$ required all past denoising steps from $T$ till $(t+1)$, this approach allows skipping some steps during sampling. More precisely, the backward diffusion process is defined as a combination of predictions of image $x_{0}$, next denoising step $x_{t}$, and random noise (with $\sigma_{t}=\eta\sqrt{\beta_{t}(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})}$ and $z_{t}\sim\mathcal{N}(0,\mathcal{I})$):
@@
While setting $\eta=1$ makes eq. 7 equivalent to eq. 6, leading to a Markovian probabilistic diffusion model, setting $\eta=0$ removes the random component from the equation, making it a Denoising Diffusion Implicit Model (DDIM), which prominent ability is to perform a deterministic mapping from given noise ($x_{T}$) to image ($x_{0}$). One of the potential benefits of implicit models is the possibility of reversing the backward diffusion process to transfer images back to the original noise. Operating in such a space by modifying resulting inversions unlocks numerous image manipulation capabilities, i.a., image editing [14, 38, 19, 41, 46, 37, 2, 54, 18, 55, 35, 12, 40, 9], image interpolation [63, 60, 47, 8], or stroke-to-image synthesis [36, 46]. The inversion process can be derived from eq. 7, leading to the formula for transferring a less noisy image $x_{t-1}$ to a more noisy one $x_{t}$:
@@
Unfortunately, a perfect image-to-noise inversion is infeasible. Due to circular dependency on $x_{t}$ within eq. 8, Dhariwal and Nichol [8] propose to approximate this equation by assuming that the model’s prediction in $t$-th step is locally equivalent to the decision in $(t-1)$-th step $\epsilon_{\theta}^{(t)}(x_{t})\approx\epsilon_{\theta}^{(t)}(x_{t-1})$. The inverted trajectory is determined in multiple steps. Hence, the error propagates further away from the image, leading to the latents that deviate from the starting noise. This flaw is presented in fig. 8.
@@
Appendix E Correlations vs. inversion steps
In this work, for the inversion process, we leverage the DDIM sampler with either $T=50$ or $T=100$ sampling steps. This choice aligns with prior works [18, 12, 24] in image edition domain, where the authors used from $50$ up to $200$ inversion steps as a proper balance between reconstruction quality and short algorithm runtime.
However, as described in Hong et al. [18], the naïve DDIM inversion procedure [50] can be reinterpreted as solving the forward diffusion ordinary differential equation (ODE) in reverse order (along the time axis) with Euler method. With this reformulation, the inversion is correct under the assumption that, with $dt$ being step size, noise predictions in $t$ and $t+dt$ steps are almost exact, thus works only when performing with many iterations.
@@
Noise ${\mathbf{x_{T}}}$ (baseline) $0.039_{\pm.00}$ $0.039_{\pm.00}$ $0.039_{\pm.00}$
@@
DDIM Latent $\mathbf{\hat{x}_{T=250}}$ $0.108_{\pm.01}$ $0.078_{\pm.01}$ $0.366_{\pm.02}$
DDIM Latent $\mathbf{\hat{x}_{T=500}}$ $0.100_{\pm.01}$ $0.069_{\pm.01}$ $0.294_{\pm.02}$
DDIM Latent $\mathbf{\hat{x}_{T=1000}}$ $0.095_{\pm.01}$ $0.065_{\pm.01}$ $0.249_{\pm.02}$
Table 11: Latent encodings resulting from the DDIM Inversion exhibit correlations, even when the procedure is performed with a lot of steps. By dividing latent encodings into $8\times 8$ patches and calculating inside the mean of top-$20$ absolute Pearson correlation coefficients, we show that DDIM latents are correlated substantially higher than Gaussian noise, even when using $T=1000$ inversion steps.
Figure 9: Approximation errors in DDIM inversion are significantly higher for plain image surfaces than for the rest of the image. Even when using $T=1000$ steps, we observe image structures in resulting latents, which can be attributed to less diverse outputs for uniform image regions.
Appendix F Most probable triangles
In section 3.2, we analyze where the latent encodings are distributed in relation to the initial noise and generated samples. In this section, we describe the methodology for obtaining such triangles, as well as investigate how those relations change depending on a number of diffusion steps $T$.
F.1 Methodology
First, we determine the vectors going from each vertex to the other vertices of the noise-sample-latent (${\mathbf{x_{T}}}$-${\mathbf{x_{0}}}$-${\mathbf{\hat{x}_{T}}}$) triangle. For sample ${\mathbf{x_{0}}}$, we obtain a vector leading to noise $\overrightarrow{{\mathbf{x_{0}}}{\mathbf{x_{T}}}}={\mathbf{x_{T}}}-{\mathbf{x_{0}}}$ and to latent $\overrightarrow{{\mathbf{x_{0}}}{\mathbf{\hat{x}_{T}}}}={\mathbf{\hat{x}_{T}}}-{\mathbf{x_{0}}}$, and calculate the angle between them using cosine similarity as

### v4

@@ 5 Conclusions
This work demonstrates that DDIM inversion errors cause latent representations to systematically deviate from a Gaussian distribution, particularly in smooth regions of the source image. We trace this to high inversion error and insufficiently diverse noise during the early noising steps, and demonstrate that this divergence degrades the quality of image editing and interpolation. We propose a simple fix by replacing initial inversion steps with forward diffusion, which successfully decorrelates the latents and significantly improves sample quality in practical applications.
@@
In the Appendix we first outline the limitations (A) of our experiments, discuss the broader impact (B) of the proposed fix, usage of LLMs (C), and list the computational resources (D) used. We then describe, in detail, the DDIM approximation error (E) and the consequences it has across varying numbers of sampling steps (F). Next, we describe our experiments: how we condition DMs (G), measuring the noise–image–latent triangles (H), identifying plain-regions pixels (I) in the image, and computing the inversion error (J). In section K, we provide the full pseudocode for the proposed fix. We demonstrate how the technique enhances image diversity and prompt alignment (L) and how it enables stochastic editing of real images (M). Finally, we compare Gaussian noise and DDIM latents in their mappings to images (N) and track how these relationships evolve during DM training (O). In section P, we include additional qualitative results: when the introduced fix might fail, and how it enhances image interpolation.
@@ Appendix A Limitations
In this work, we analyzed the relation between the random noise, image generations and their latent encodings obtained through DDIM Inversion. While our studies focused on DDIM approximation error from Song et al. (2021), there exist other solvers and inversion methods, as described in Section 2, bringing their own advantages and limitations. The error of DDIM inversion strongly depends on the number of steps with which it is performed. In particular, performing the process very granularly, e.g., using $T=1000$ steps, can result in strong suppression of the correlation. Nevertheless, the default number of steps we have chosen, i.e., 100, is, according to previous works (Hong et al., 2024; Garibi et al., 2024; Kim et al., 2022b), a practical choice as a good balance between the reconstruction error and the speed of the algorithm. In Appendix F, we present that the latents exhibit correlations when using $1000$ sampling steps, and that the proposed fix can help also in such case.
@@
The observations from our analytical experiments (correlations in Table 2, interpolations in Fig. 5) generalize well to all tested diffusion models, but are less evident in the LDM model trained on the CelebA-HQ images. We attribute this exemption to the specificity of the dataset on which the model was trained - photos with centered human faces, usually with uniform backgrounds. We believe that, unlike models trained on a larger number of concepts, the process of generating faces with uniform backgrounds is more stable and introduces little detail in subsequent steps, making the difference in approximation error for plain and non-plain areas less significant.
@@
As mentioned in Section 2, the DDIM inversion error can be detrimental when using a small number of steps. Even though the solution proposed in this work (involving forward diffusion in first inversion steps) drastically removes correlations in latents and, thus, improves image interpolation and editing, it does not improve the numerical inversion error resulting from using small number of steps. In our experiments with 50 steps that are commonly used for edition, we show no significant drawbacks. However, in the extreme cases, using our fix in even a single step, might result in the loss of information necessary for correct image reconstruction, hence it may be then less preferred than standard DDIM inversion. In Section P.1, we present failure cases for introduced solution.
Appendix B LLM Usage
Throughout the preparation of this manuscript, we employed a large language model (LLM) as a writing assistant. Its use was focused on improving the clarity and readability of the text, correcting grammar, and refining sentence structure. The authors carefully reviewed, edited, and take full responsibility for all content, ensuring the scientific integrity and accuracy of the final paper.
Appendix C Broader impact
@@
As our work is mostly analytical, we do not provide new technologies that might have a significant societal impact. However, our solution for improving the accuracy of DDIM inversion has potential implications that extend beyond technical advancements in diffusion models. As our fix enables more prompt-aligned image editing it could be misused to advance image manipulation techniques. The enhanced interpolation quality could make synthetic content more convincing and harder to detect. The authors do not endorse using the method for deceptive or malicious purposes, and discourage any application that could erode trust or cause harm.
Appendix D Compute resources
@@
Almost all the experiments we perform are based on performing a sampling process using a diffusion model from noise, performing DDIM inversion and, possibly, image reconstruction from the latent, where each of these processes takes the same number of steps, hence the same number of GPU-hours on average. As our experiments differ in terms of number of sampling steps and images to generate, we provide average GPU time per one sampling step per batch (with $B$ denoting batch size) for each model as following: ADM-32 ($B=256$): 0.054s, ADM-64 ($B=128$): 0.093s, ADM-256 ($B=64$): 0.901s, LDM ($B=128$): 0.273s, DiT ($B=128$): 0.104s, IF ($B=64$): 0.609s. Note that some experiments, such as analyzing inversion approximation errors per step (Fig. 4) or sampling from noise interpolations (Fig. 5), involves performing the procedures several times. Taking into consideration all the experiments described in the main text of this work, fully reproducing them takes roughly 110 GPU hours. However, considering the prototyping time, preliminary and failed experiments, as well as the fact that most of the experiments must be performed sequentially (e.g., inversion after image generation, image reconstruction after inversion), the overall execution time of the entire research project is multiple times longer.
Appendix E Approximation error in DDIM Inversion
Denoising Diffusion Probabilistic Models (DDPMs, Ho et al. (2020)) generate samples by reversing the forward diffusion process, modeled as a Markov Chain, where a clean image $x_{0}$ is progressively transformed to white Gaussian noise $x_{T}$ in $T$ diffusion steps. A partially noised image $x_{t}$, which serves as an intermediate object in this process, is expressed as $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$ where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, $\alpha_{t}=1-\beta_{t}$, $\epsilon_{t}\sim\mathcal{N}(0,\mathcal{I})$, and $\{\beta_{t}\}_{t=1}^{T}$ is a variance schedule, controlling how much of the noise is contained in the image at the specific step $t$.
@@ Appendix E Approximation error in DDIM Inversion
Song et al. (2021) reformulate the diffusion process as a non-Markovian, which leads to a speed-up of the sampling process. While previously obtaining a less noisy image at $x_{t}$ required all past denoising steps from $T$ till $(t+1)$, this approach allows skipping some steps during sampling. More precisely, the backward diffusion process is defined as a combination of predictions of image $x_{0}$, next denoising step $x_{t}$, and random noise (with $\sigma_{t}=\eta\sqrt{\beta_{t}(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})}$ and $z_{t}\sim\mathcal{N}(0,\mathcal{I})$):
@@
While setting $\eta=1$ makes Eq. 7 equivalent to Eq. 6, leading to a Markovian probabilistic diffusion model, setting $\eta=0$ removes the random component from the equation, making it a Denoising Diffusion Implicit Model (DDIM), which prominent ability is to perform a deterministic mapping from given noise ($x_{T}$) to image ($x_{0}$). One of the potential benefits of implicit models is the possibility of reversing the backward diffusion process to transfer images back to the original noise. Operating in such a space by modifying resulting inversions unlocks numerous image manipulation capabilities, i.a., image editing (Hertz et al., 2022; Mokady et al., 2023; Huberman-Spiegelglas et al., 2024; Parmar et al., 2023; Rout et al., 2025; Miyake et al., 2023; Brack et al., 2024; Tang et al., 2024; Hong et al., 2024; Wallace et al., 2023; Samuel et al., ; Garibi et al., 2024; Pan et al., 2023; Dong et al., 2023), image interpolation (Zheng et al., 2024; Zhang et al., 2024b; Samuel et al., 2023; Dhariwal & Nichol, 2021), or stroke-to-image synthesis (Meng et al., 2022; Rout et al., 2025). The inversion process can be derived from Eq. 7, leading to the formula for transferring a less noisy image $x_{t-1}$ to a more noisy one $x_{t}$:
@@
Unfortunately, a perfect image-to-noise inversion is infeasible. Due to circular dependency on $x_{t}$ within Eq. 8, Dhariwal & Nichol (2021) propose to approximate this equation by assuming that the model’s prediction in $t$-th step is locally equivalent to the decision in (t1)(t-1)-th step $\epsilon_{\theta}^{(t)}(x_{t})\approx\epsilon_{\theta}^{(t)}(x_{t-1})$. The inverted trajectory is determined in multiple steps. Hence, the error propagates further away from the image, leading to the latents that deviate from the starting noise. This flaw is presented in Fig. 8.
@@
Appendix F Correlations vs. inversion steps
In this work, for the inversion process, we leverage the DDIM sampler with either $T=50$ or $T=100$ sampling steps. This choice aligns with prior works (Hong et al., 2024; Garibi et al., 2024; Kim et al., 2022b) in image edition domain, where the authors used from $50$ up to $200$ inversion steps as a proper balance between reconstruction quality and short algorithm runtime.
However, as described in Hong et al. (2024), the naïve DDIM inversion procedure (Song et al., 2021) can be reinterpreted as solving the forward diffusion ordinary differential equation (ODE) in reverse order (along the time axis) with Euler method. With this reformulation, the inversion is correct under the assumption that, with $dt$ being step size, noise predictions in $t$ and $t+dt$ steps are almost exact, thus works only when performing with many iterations.
@@
For a more thorough analysis, in Table 9 we evaluate how our fix decorraltes latents in situation where we use $T=1000$ inversion steps. As visible, replacing just $1$ step of DDIM Inversion with forward diffusion significantly reduces correlation at minimal loss in image reconstruction.
@@
Noise $\mathbf{x_{T}}$} (baseline) $039_{\pm.00} 0.0$$39_{\pm.00} 0.0$$039_{\pm.00}$
@@
DDIM Latent $\mathbf{\hat{x}_{T=250}}$ $108_{\pm.01} 0.0$$078_{\pm.01}$ $0.366_{\pm.02}$
DDIM Latent $\mathbf{\hat{x}_{T=500}}$ $100_{\pm.01} 0.0$$069_{\pm.01}$ $0.294_{\pm.02}$
DDIM Latent $\mathbf{\hat{x}_{T=1000}}$ $095_{\pm.01} 0.0$$065_{\pm.01}$ $0.249_{\pm.02}$
Table 8: Latent encodings resulting from the DDIM Inversion exhibit correlations, even when the procedure is performed with a lot of steps. By dividing latent encodings into $8\times 8$ patches and calculating the mean of top-$20$ Pearson coefficients, we show that DDIM latents are correlated substantially higher than Gaussian noise, even when using $T=1000$ inversion steps.
Model Prior Corr. $\downarrow$ KL $\times 10^{-4}\downarrow$ Image Recon. $\downarrow$
Noise (upper bound) $0.039_{\pm.00}$ 4.832 0.000
DDIM Latent $0.095_{\pm.01}$ 47.015 0.014
w/ our fix (4%) $0.055_{\pm.00}$ 5.125 0.033
ADM-64 w/ our fix (2%) $0.055_{\pm.00}$ 5.363 0.024
(T=1000) w/ our fix (1%) $0.055_{\pm.00}$ 5.679 0.019
w/ our fix (0.5%) $0.055_{\pm.00}$ 6.057 0.017
w/ our fix (0.2%) $0.055_{\pm.00}$ 6.662 0.015
w/ our fix (0.1%) $0.055_{\pm.00}$ 7.228 0.014
Noise (upper bound) $0.039_{\pm.00}$ 4.832 0.000
DDIM Latent $0.065_{\pm.01}$ 17.931 0.009
w/ our fix (4%) $0.057_{\pm.00}$ 5.233 0.060
DiT w/ our fix (2%) $0.057_{\pm.00}$ 5.071 0.041
(T=1000) w/ our fix (1%) $0.057_{\pm.00}$ 5.850 0.029
w/ our fix (0.5%) $0.058_{\pm.00}$ 7.029 0.021
w/ our fix (0.2%) $0.058_{\pm.00}$ 8.045 0.016
w/ our fix (0.1%) $0.058_{\pm.00}$ 8.409 0.014
Noise (upper bound) $0.039_{\pm.00}$ 4.832 0.000
DDIM Latent $0.249_{\pm.02}$ 63.962 0.044
w/ our fix (4%) $0.055_{\pm.00}$ 5.024 0.037
IF w/ our fix (2%) $0.055_{\pm.00}$ 5.215 0.031
(T=1000) w/ our fix (1%) $0.055_{\pm.00}$ 5.706 0.027
w/ our fix (0.5%) $0.055_{\pm.00}$ 6.117 0.026
w/ our fix (0.2%) $0.055_{\pm.00}$ 5.562 0.025
w/ our fix (0.1%) $0.056_{\pm.00}$ 5.349 0.024
Table 9: Latent correlations, KL divergence to random Gaussian noise, and image reconstruction error across models (ADM-64, DiT, IF) when using $T=1000$ inversion steps. We show that using our simple fix in just one step of inversion process significantly decorellates DDIM latents with minimal loss in image reconstruction performance.
Figure 9: Approximation errors in DDIM inversion are significantly higher for plain image surfaces than for the rest of the image. Even when using $T=1000$ steps, we observe image structures in DDIM latents, notably for uniform image regions.
Appendix G Conditioning Diffusion Models
For a thorough analysis, we employ both unconditional (ADM-32, ADM-64, ADM-256) and conditional (DiT, IF, SDXL) diffusion models. For conditioning, we take prompts from the Recap-COCO-30K dataset (Li et al., 2025) for IF and Stable Diffusion XL and ImageNet class names for DiT. However, as noted by Mokady et al. (2023), Classifier-Free Guidance introduces additional errors to the DDIM inversion, so to focus solely on the inversion approximation error, we disable CFG by setting the guidance scale to $w=1$.
Appendix H Most probable triangles
In Section 3.2, we analyze where the latent encodings are distributed in relation to the initial noise and generated samples. Here, we determine the most probable angles formed by noises $\mathbf{x_{T}}$}, samples $\mathbf{x_{0}}$}, and latents $\mathbf{\hat{x}_{T}}$} for each model.
H.1 Methodology
First, we determine the vectors going from each vertex to the other vertices of the noise-sample-latent ($\mathbf{x_{T}}$$\mathbf{x_{0}}$$\mathbf{\hat{x}_{T}}$}) triangle. For sample $\mathbf{x_{0}}$}, we obtain a vector leading to noise $\overrightarrow{{\mathbf{x_{0}}}{\mathbf{x_{T}}}}={\mathbf{x_{T}}}-{\mathbf{x_{0}}}$ and to latent $\overrightarrow{{\mathbf{x_{0}}}{\mathbf{\hat{x}_{T}}}}={\mathbf{\hat{x}_{T}}}-{\mathbf{x_{0}}}$, and calculate the angle between them using cosine similarity as